Jack is 4 times as old as Lacy. 9 years from now the sum of their ages will be 63. How old are they now?

Answer by addingup(3677)

J = 4L
J+9+L+9 = 63 Substitute for J and add on left:
4L+L+18 = 63 Add L on left and subtract 18 on both sides:
5L = 45
L = 9
Lacy is 9 and Jack is 9*4 = 36
In 9 years:
9+9 = 18 Lacy
36+9= 45 Jack
. . ------
Total 63 Correct
John